SPECTROSCOPIC CHARACTERIZATION OF HOONO AND ITS BINDING ENERGY VIA INFRARED ACTION SPECTROSCOPY

\underline{ILANA B. POLLACK}, EUNICE X.J. LI, IAN M. KONEN AND MARSHA I. LESTER, Department of Chemistry, University of Pennsylvania, Philadelphia, PA 19104-6323.

The OH + NO_{2} + M → HONO_{2} + M reaction is of fundamental importance in atmospheric chemistry because it is an important sink of reactive HO_{x} and NO_{x} radicals that directly affect the destruction of ozone in the stratosphere and the production of ozone in the troposphere. Peroxynitrous acid (HOONO), a secondary product of the OH + NO_{2} reaction and a less stable isomer of HONO_{2}, has been created in a pulsed supersonic free-jet expansion by reaction of photolytically generated OH radicals with NO_{2}. An infrared pump-ultraviolet probe technique has been used to spectroscopically characterize the trans-perp (tp) conformer of HOONO. Infrared action spectra of tp-HOONO have been recorded in the OH overtone region near 1.4 μm using a single-mode optical parametric oscillator. Rotational structure of the pure OH overtone band has been observed at 6971.4 cm^{−1} (band origin). Experimental spectra have been assigned by comparison to simulated rotational band structure using rotational constants and estimated vibrational frequencies predicted by previous ab initio calculations. Infrared excitation of tp-HOONO in the OH overtone region provides sufficient energy to break the O–O bond leading to unimolecular dissociation into OH and NO_{2} products in less than 6 ns (laser-limited). An OH (v=0) product state distribution was obtained following infrared overtone excitation. The highest populated OH 2Π_{3/2} (v=0) state observed, J = 15/2, establishes a tp-HOONO binding energy of D_{0} ≤ 17 kcal/mol, which is consistent with current theoretical estimates of the binding energy.